Bus Econ 1.2.37 Question Help manager of a restaurant found that the cost to produce 100 cups of coffee is $10.74, while the cost to produce 400 cups is $38.94. Assume the cost Clx) isa inear function of x, the number of cups produced. Answer parts a through f for C(x). C(x)(Use integers or decimals for any numbers in the expression.)
Answer #1

Marginal cost of a cup of coffee=change in Cost/change in cups of coffee

marginal cost=28.20/300=0.094

and variable cost for initial 100 cups=9.4

thus TFC=10.74-9.4=1.34

Thus total cost=TFC+TVC

TC(c)=1.34+0.094x